# Dynamics of the general $Q$-tensor model interacting with a rigid body

### v1

@@ 3. Energy decay and equilibria of the interaction problem
In this section, we show that the total energy of the interaction problem (2.5)–(2.9) is a non-increasing functional, and we deduce the set of equilibria. First, we decompose $\tau=\tau_{\mathrm{h}}+\tau_{\mathrm{l}}$ and $H=H_{\mathrm{h}}+H_{\mathrm{l}}$ into the respective higher and lower order terms. Invoking the commutator $[A,B]=AB-BA$ and anticommutator $\{A,B\}=AB+BA$ of $A$, $B\in\mathbb{C}^{3\times 3}$, and writing $\mathbb{D}=\mathbb{D}(u^{\mathcal{F}})$ and $\mathbb{W}=\mathbb{W}(u^{\mathcal{F}})$, we get
@@
$H_{\mathrm{h}}(Q^{\mathcal{F}})$ $=\Delta Q^{\mathcal{F}}-Q^{\mathcal{F}},\enspace H_{\mathrm{l}}(Q^{\mathcal{F}})=\bigl{(}(Q^{\mathcal{F}})^{2}-\operatorname{tr}\bigl{(}(Q^{\mathcal{F}})^{2}\bigr{)}\nicefrac{{\operatorname{Id}_{3}}}{{3}}\bigr{)}-\operatorname{tr}\bigl{(}(Q^{\mathcal{F}})^{2}\bigr{)}Q^{\mathcal{F}},$
@@
$\tau_{\mathrm{h}}(Q^{\mathcal{F}})$ $=-\xi\bigl{(}\nicefrac{{2}}{{3}}H_{\mathrm{h}}+\{Q^{\mathcal{F}},H_{\mathrm{h}}\}-2(Q^{\mathcal{F}}+\nicefrac{{\operatorname{Id}_{3}}}{{3}})\operatorname{tr}(Q^{\mathcal{F}}H_{\mathrm{h}})\bigr{)}\enspace\text{and}\enspace$
@@
(3.2) $S_{\xi}(Q^{\mathcal{F}})A$ $\coloneqq[Q^{\mathcal{F}},A]-\nicefrac{{2\xi}}{{3}}A-\xi\{Q^{\mathcal{F}},A\}+2\xi(Q^{\mathcal{F}}+\nicefrac{{\operatorname{Id}_{3}}}{{3}})\operatorname{tr}(Q^{\mathcal{F}}A)\enspace\text{and}\enspace$
@@
For real $u^{\mathcal{F}}$, $Q^{\mathcal{F}}$, we get $\tau_{\mathrm{h}}(Q^{\mathcal{F}})+\sigma(Q^{\mathcal{F}})=S_{\xi}(Q^{\mathcal{F}})(\Delta-\operatorname{Id}_{3})Q^{\mathcal{F}}$ and $-S(\nabla u^{\mathcal{F}},Q^{\mathcal{F}})=\tilde{S}_{\xi}(Q^{\mathcal{F}})\nabla u^{\mathcal{F}}$. The following lemma, for which we refer to [21, Lemma 5.1], establishes a link between $S_{\xi}$ and $\tilde{S}_{\xi}$. Note that we also denote the product $\langle A,B\rangle_{\mathbb{C}^{3\times 3}}$ by $A:B$, and it is defined by $\operatorname{tr}(A\overline{B})$.
[…]
@@
(3.3) $\mathrm{E}(t)\coloneqq\mathrm{E}_{\mathrm{kin}}(t)+\mathrm{E}_{Q}(t)+\mathrm{E}_{\mathrm{trans}}(t)+\mathrm{E}_{\mathrm{rot}}(t),\enspace\text{where}\enspace$
$\mathrm{E}_{\mathrm{kin}}(t)=\frac{1}{2}\int_{\mathcal{F}(t)}|u^{\mathcal{F}}(t,x)|^{2}\,\mathrm{d}x,\enspace\mathrm{E}_{\mathrm{trans}}(t)=\frac{1}{2}m_{\mathcal{S}}|h^{\prime}(t)|^{2}\enspace\text{and}\enspace\mathrm{E}_{\mathrm{rot}}(t)=\frac{1}{2}(J\Omega\cdot\Omega)(t).$
@@
[…]
Let $p>4$ and $z_{0}=(v_{0},Q_{0},\ell_{0},\omega_{0})\in\mathrm{X}_{\gamma}$, for $\mathrm{X}_{\gamma}$ as introduced in (2.10), recall $H$ from (2.3), and consider a solution $(u^{\mathcal{F}},Q^{\mathcal{F}},h^{\prime},\Omega,\pi^{\mathcal{F}})$ to (2.5)–(2.9) in the regularity class
$\mathrm{W}^{1,p}\bigl{(}0,T;\mathrm{L}^{2}(\mathcal{F}(\cdot))^{3}\bigr{)}\cap\mathrm{L}^{p}\bigl{(}0,T;\mathrm{H}^{2}(\mathcal{F}(\cdot))^{3}\bigr{)}\times\mathrm{W}^{1,p}\bigl{(}0,T^{\prime};\mathrm{H}^{1}(\mathcal{F}(\cdot),\mathbb{S}_{0,\mathbb{R}}^{3})\bigr{)}\cap\mathrm{L}^{p}\bigl{(}0,T^{\prime};\mathrm{H}^{3}(\mathcal{F}(\cdot),\mathbb{S}_{0,\mathbb{R}}^{3})\bigr{)}$
×$\times\mathrm{W}^{1,p}(0,T)^{3}\times\mathrm{W}^{1,p}(0,T)^{3}\times\mathrm{L}^{p}\bigl{(}0,T;\mathrm{H}^{1}(\mathcal{F}(\cdot))\cap\mathrm{L}_{0}^{2}(\mathcal{F}(\cdot))\bigr{)}.$
@@
    We have =$\mathcal{E}=\{(0,Q_{*},0,0):H(Q_{*})=0,\text{ in }\mathcal{F}_{0},\enspace\partial_{\nu}Q_{*}=0,\text{ on }\partial\mathcal{F}_{0}\}$ for the set of equilibria.
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{E}_{\mathrm{kin}}(t)=-\int_{\mathcal{F}(t)}|\nabla u^{\mathcal{F}}|^{2}\,\mathrm{d}x-\int_{\mathcal{F}(t)}(\tau(Q^{\mathcal{F}})+\sigma(Q^{\mathcal{F}})):\nabla u^{\mathcal{F}}\,\mathrm{d}x-\int_{\partial\mathcal{S}(t)}\tilde{\Sigma}(u^{\mathcal{F}},\pi^{\mathcal{F}},Q^{\mathcal{F}})n\cdot u^{\mathcal{F}}\,\mathrm{d}\Gamma.$
In the same way as in [9, (3.15) and (3.17)], we find $\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{E}_{\mathrm{trans}}(t)+\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{E}_{\mathrm{rot}}(t)=\int_{\partial\mathcal{S}(t)}\tilde{\Sigma}(u^{\mathcal{F}},\pi^{\mathcal{F}},Q^{\mathcal{F}})n\cdot u^{\mathcal{F}}\,\mathrm{d}\Gamma$. A concatenation of the latter two identities leads to
@@
It remains to handle $\mathrm{E}_{Q}$ from (2.1). For this purpose, we multiply the equation (2.5)2 satisfied by the molecular orientation $Q^{\mathcal{F}}$ from the right by $-H$ and integrate over $\mathcal{F}(t)$. This yields
(3.5) $-\int_{\mathcal{F}(t)}\partial_{t}Q^{\mathcal{F}}:H\,\mathrm{d}x-\int_{\mathcal{F}(t)}(u^{\mathcal{F}}\cdot\nabla)Q^{\mathcal{F}}:H\,\mathrm{d}x+\int_{\mathcal{F}(t)}S(\nabla u^{\mathcal{F}},Q^{\mathcal{F}}):H\,\mathrm{d}x=-\int_{\mathcal{F}(t)}\operatorname{tr}(H^{2})\,\mathrm{d}x.$
From Section 3 and the aforementioned identities $\tau_{\mathrm{h}}(Q^{\mathcal{F}})+\sigma(Q^{\mathcal{F}})=S_{\xi}(Q^{\mathcal{F}})((\Delta-\operatorname{Id}_{3})Q^{\mathcal{F}})$ as well as $-S(\nabla u^{\mathcal{F}},Q^{\mathcal{F}})=\tilde{S}_{\xi}(Q^{\mathcal{F}})\nabla u^{\mathcal{F}}$ and $S_{\xi}(Q^{\mathcal{F}})H_{\mathrm{l}}=\tau_{\mathrm{l}}(Q^{\mathcal{F}})+\nabla Q^{\mathcal{F}}\odot\nabla Q^{\mathcal{F}}$, we deduce that
@@
$-\int_{\mathcal{F}(t)}\partial_{t}Q^{\mathcal{F}}:H\,\mathrm{d}x+\int_{\partial\mathcal{S}(t)}\Bigl{(}\frac{1}{2}|\nabla Q^{\mathcal{F}}|^{2}+\frac{1}{2}\operatorname{tr}\bigl{(}(Q^{\mathcal{F}})^{2}\bigr{)}-\frac{1}{3}\operatorname{tr}\bigl{(}(Q^{\mathcal{F}})^{3}\bigr{)}+\frac{1}{4}\operatorname{tr}\bigl{(}(Q^{\mathcal{F}})^{2}\bigr{)}^{2}\Bigr{)}(u^{\mathcal{F}}\cdot n)\,\mathrm{d}\Gamma.$
@@
$\int_{\partial\mathcal{S}(t)}\Bigl{(}\frac{1}{2}|\nabla Q^{\mathcal{F}}|^{2}+\frac{1}{2}\operatorname{tr}\bigl{(}(Q^{\mathcal{F}})^{2}\bigr{)}-\frac{1}{3}\operatorname{tr}\bigl{(}(Q^{\mathcal{F}})^{3}\bigr{)}+\frac{1}{4}\operatorname{tr}\bigl{(}(Q^{\mathcal{F}})^{2}\bigr{)}^{2}\Bigr{)}(u^{\mathcal{F}}\cdot n)\,\mathrm{d}\Gamma.$
For verifying (a), it thus remains to show that the last three terms in (3.7) cancel out. To this end, we address $\int_{\mathcal{F}(t)}(u^{\mathcal{F}}\cdot\nabla)Q^{\mathcal{F}}:H\,\mathrm{d}x$ and discuss the resulting integrals for each term in $H$. Integrating by parts, and employing $\operatorname{div}u^{\mathcal{F}}=0$, (2.8) and the identities $(u^{\mathcal{F}}\cdot\nabla)((Q^{\mathcal{F}})^{2})=(u^{\mathcal{F}}\cdot\nabla)Q^{\mathcal{F}}\cdot Q^{\mathcal{F}}+Q^{\mathcal{F}}\cdot(u^{\mathcal{F}}\cdot\nabla)Q^{\mathcal{F}}$, $-\int_{\mathcal{F}(t)}(u^{\mathcal{F}}\cdot\nabla)Q^{\mathcal{F}}:(\operatorname{tr}((Q^{\mathcal{F}})^{2})\nicefrac{{\operatorname{Id}_{3}}}{{3}})\,\mathrm{d}x=0$ thanks to $\operatorname{tr}(Q^{\mathcal{F}})=0$ as well as the identity involving the above inner product $Q^{\mathcal{F}}:\bigl{(}(u^{\mathcal{F}}\cdot\nabla)(\operatorname{tr}((Q^{\mathcal{F}})^{2})Q^{\mathcal{F}})\bigr{)}=3(u^{\mathcal{F}}\cdot\nabla)Q^{\mathcal{F}}:\bigl{(}\operatorname{tr}((Q^{\mathcal{F}})^{2})Q^{\mathcal{F}}\bigr{)}$, we find that
@@
$=-\int_{\partial\mathcal{S}(t)}\Bigl{(}\frac{1}{2}\operatorname{tr}\bigl{(}(Q^{\mathcal{F}})^{2}\bigr{)}-\frac{1}{3}\operatorname{tr}\bigl{(}(Q^{\mathcal{F}})^{3}\bigr{)}+\frac{1}{4}\operatorname{tr}\bigl{(}(Q^{\mathcal{F}})^{2}\bigr{)}^{2}\Bigr{)}(u^{\mathcal{F}}\cdot n)\,\mathrm{d}\Gamma.$
@@
For the second term, we obtain $-\sum_{i,j,k=1}^{3}\int_{\mathcal{F}(t)}u_{k}^{\mathcal{F}}\partial_{k}\nabla Q_{ij}^{\mathcal{F}}\cdot\nabla Q_{ji}^{\mathcal{F}}\,\mathrm{d}x=-\frac{1}{2}\int_{\partial\mathcal{S}(t)}|\nabla Q^{\mathcal{F}}|^{2}(u^{\mathcal{F}}\cdot n)\,\mathrm{d}\Gamma$, so
(3.9) $\int_{\mathcal{F}(t)}(u^{\mathcal{F}}\cdot\nabla)Q^{\mathcal{F}}:\Delta Q^{\mathcal{F}}\,\mathrm{d}x=-\int_{\mathcal{F}(t)}\nabla u^{\mathcal{F}}:(\nabla Q^{\mathcal{F}}\odot\nabla Q^{\mathcal{F}})\,\mathrm{d}x-\frac{1}{2}\int_{\partial\mathcal{S}(t)}|\nabla Q^{\mathcal{F}}|^{2}(u^{\mathcal{F}}\cdot n)\,\mathrm{d}\Gamma.$
@@
Concerning (b), we argue that the equilibria $\mathcal{E}$ are the critical points of the energy functional $\mathrm{E}$, so assume $\left.\frac{\,\mathrm{d}}{\,\mathrm{d}t}\mathrm{E}(t)\right|_{t=t_{0}}=0$ at $t_{0}>0$. From (a), it follows that $\nabla u^{\mathcal{F}}=0$ and thus also $u^{\mathcal{F}}=0$ in $\mathcal{F}(t_{0})$ thanks to $u^{\mathcal{F}}=0$ on $\partial\mathcal{O}$. As in the proof of [35, Lemma 2.1], we deduce that $h^{\prime}=\Omega=0$. The relation in (a) also shows that $H(Q_{*})=0$ has to hold for an equilibrium, so $\mathcal{E}$ is of the asserted shape. ∎
[…]
Combining Section 3(b) and [21, Remark 3.3], stating that $H(Q_{*})=0$ implies $Q_{*}=0$ for spatially constant $Q_{*}$, we conclude that $(0,0,0,0)$ is the only such equilibrium to (2.5)–(2.8).

### v2

@@ 3. Energy decay and equilibria of the interaction problem
In this section, we show that the total energy of the interaction problem (2.5)–(2.9) is a non-increasing functional, and we deduce the set of equilibria. First, we decompose $\tau=\tau_{\mathrm{h}}+\tau_{\mathrm{l}}$ and $H=H_{\mathrm{h}}+H_{\mathrm{l}}$ into the respective higher and lower order terms. Invoking the commutator [A,B]=ABBA[A,B]=AB-BA and anticommutator $\{A,B\}=AB+BA$ of $A$, $B\in\mathbb{C}^{3\times 3}$, and writing $\mathbb{D}=\mathbb{D}(u^{\mathcal{F}})$ and $\mathbb{W}=\mathbb{W}(u^{\mathcal{F}})$, we get
@@
$H_{\mathrm{h}}(Q^{\mathcal{F}})$ $=\Delta Q^{\mathcal{F}}-Q^{\mathcal{F}},\kern 5.0ptH_{\mathrm{l}}(Q^{\mathcal{F}})=\bigl((Q^{\mathcal{F}})^{2}-\operatorname{tr}\bigl((Q^{\mathcal{F}})^{2}\bigr)\nicefrac{{\operatorname{Id}_{3}}}{{3}}\bigr)-\operatorname{tr}\bigl((Q^{\mathcal{F}})^{2}\bigr)Q^{\mathcal{F}},$
@@
$\tau_{\mathrm{h}}(Q^{\mathcal{F}})$ $=-\xi\bigl(\nicefrac{{2}}{{3}}H_{\mathrm{h}}+\{Q^{\mathcal{F}},H_{\mathrm{h}}\}-2(Q^{\mathcal{F}}+\nicefrac{{\operatorname{Id}_{3}}}{{3}})\operatorname{tr}(Q^{\mathcal{F}}H_{\mathrm{h}})\bigr)\kern 5.0pt\text{and}\kern 5.0pt$
@@
(3.2) $S_{\xi}(Q^{\mathcal{F}})A$ $\coloneqq[Q^{\mathcal{F}},A]-\nicefrac{{2\xi}}{{3}}A-\xi\{Q^{\mathcal{F}},A\}+2\xi(Q^{\mathcal{F}}+\nicefrac{{\operatorname{Id}_{3}}}{{3}})\operatorname{tr}(Q^{\mathcal{F}}A)\kern 5.0pt\text{and}\kern 5.0pt$
@@
For real $u^{\mathcal{F}}$, $Q^{\mathcal{F}}$, we get $\tau_{\mathrm{h}}(Q^{\mathcal{F}})+\sigma(Q^{\mathcal{F}})=S_{\xi}(Q^{\mathcal{F}})(\Delta-\operatorname{Id}_{3})Q^{\mathcal{F}}$ and $S(\nabla u^{\mathcal{F}},Q^{\mathcal{F}})=\tilde{S}_{\xi}(Q^{\mathcal{F}})\nabla u^{\mathcal{F}}$. The following lemma, for which we refer to [22, Lemma 5.1], establishes a link between $S_{\xi}$ and $\tilde{S}_{\xi}$. Note that we also denote the product $\langle A,B\rangle_{\mathbb{C}^{3\times 3}}$ by $A:B$, and it is defined by $\operatorname{tr}(A\overline{B})$.
Lemma 3.1.
@@
(3.3) $\mathrm{E}(t)\coloneqq\mathrm{E}_{\mathrm{kin}}(t)+\mathrm{E}_{Q}(t)+\mathrm{E}_{\mathrm{trans}}(t)+\mathrm{E}_{\mathrm{rot}}(t),\kern 5.0pt\text{where}\kern 5.0pt$
$\mathrm{E}_{\mathrm{kin}}(t)=\frac{1}{2}\int_{\mathcal{F}(t)}|u^{\mathcal{F}}(t,x)|^{2}\,\mathrm{d}x,\kern 5.0pt\mathrm{E}_{\mathrm{trans}}(t)=\frac{1}{2}m_{\mathcal{S}}|h^{\prime}(t)|^{2}\kern 5.0pt\text{and}\kern 5.0pt\mathrm{E}_{\mathrm{rot}}(t)=\frac{1}{2}(J\Omega\cdot\Omega)(t).$
@@
Proposition 3.2.
Let $p>4$ and $z_{0}=(v_{0},Q_{0},\ell_{0},\omega_{0})\in\mathrm{X}_{\gamma}$, for $\mathrm{X}_{\gamma}$ as introduced in (2.10), recall $H$ from (2.3), and consider a solution $(u^{\mathcal{F}},Q^{\mathcal{F}},h^{\prime},\Omega,\pi^{\mathcal{F}})$, whose existence will be discussed in Theorem 4.1, to (2.5)–(2.9) in the regularity class
$\mathrm{W}^{1,p}\bigl(0,T;\mathrm{L}^{2}(\mathcal{F}(\cdot))^{3}\bigr)\cap\mathrm{L}^{p}\bigl(0,T;\mathrm{H}^{2}(\mathcal{F}(\cdot))^{3}\bigr)\times\mathrm{W}^{1,p}\bigl(0,T;\mathrm{H}^{1}(\mathcal{F}(\cdot),\mathbb{S}_{0,\mathbb{R}}^{3})\bigr)\cap\mathrm{L}^{p}\bigl(0,T;\mathrm{H}^{3}(\mathcal{F}(\cdot),\mathbb{S}_{0,\mathbb{R}}^{3})\bigr)$
$\times\mathrm{W}^{1,p}(0,T)^{3}\times\mathrm{W}^{1,p}(0,T)^{3}\times\mathrm{L}^{p}\bigl(0,T;\mathrm{H}^{1}(\mathcal{F}(\cdot))\cap\mathrm{L}_{0}^{2}(\mathcal{F}(\cdot))\bigr).$
@@
    We have $\mathcal{E}=\{(0,Q_{*},0,0):H(Q_{*})=0,\text{ in }\mathcal{F}_{0},\kern 5.0pt\partial_{\nu}Q_{*}=0,\text{ on }\partial\mathcal{F}_{0}\}$ for the set of equilibria.
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{E}_{\mathrm{kin}}(t)=-\int_{\mathcal{F}(t)}|\nabla u^{\mathcal{F}}|^{2}\,\mathrm{d}x-\int_{\mathcal{F}(t)}(\tau(Q^{\mathcal{F}})+\sigma(Q^{\mathcal{F}})):\nabla u^{\mathcal{F}}\,\mathrm{d}x-\int_{\partial\mathcal{S}(t)}\tilde{\Sigma}(u^{\mathcal{F}},\pi^{\mathcal{F}},Q^{\mathcal{F}})\nu(t)\cdot u^{\mathcal{F}}\,\mathrm{d}\Gamma.$
Similarly as in [9, (3.15) and (3.17)], we find $\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{E}_{\mathrm{trans}}(t)+\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{E}_{\mathrm{rot}}(t)=\int_{\partial\mathcal{S}(t)}\tilde{\Sigma}(u^{\mathcal{F}},\pi^{\mathcal{F}},Q^{\mathcal{F}})\nu(t)\cdot u^{\mathcal{F}}\,\mathrm{d}\Gamma$. A concatenation of the latter two identities leads to
@@
It remains to handle $\mathrm{E}_{Q}$ from (2.1). For this purpose, we multiply the equation (2.5)2 satisfied by the molecular orientation $Q^{\mathcal{F}}$ from the right by H-H and integrate over $\mathcal{F}(t)$. This yields
(3.5) $\int_{\mathcal{F}(t)}\partial_{t}Q^{\mathcal{F}}:H\,\mathrm{d}x-\int_{\mathcal{F}(t)}(u^{\mathcal{F}}\cdot\nabla)Q^{\mathcal{F}}:H\,\mathrm{d}x+\int_{\mathcal{F}(t)}S(\nabla u^{\mathcal{F}},Q^{\mathcal{F}}):H\,\mathrm{d}x=-\int_{\mathcal{F}(t)}\operatorname{tr}(H^{2})\,\mathrm{d}x.$
From Section 3 and the aforementioned identities $\tau_{\mathrm{h}}(Q^{\mathcal{F}})+\sigma(Q^{\mathcal{F}})=S_{\xi}(Q^{\mathcal{F}})((\Delta-\operatorname{Id}_{3})Q^{\mathcal{F}})$ as well as $S(\nabla u^{\mathcal{F}},Q^{\mathcal{F}})=\tilde{S}_{\xi}(Q^{\mathcal{F}})\nabla u^{\mathcal{F}}$ and $S_{\xi}(Q^{\mathcal{F}})H_{\mathrm{l}}=\tau_{\mathrm{l}}(Q^{\mathcal{F}})+\nabla Q^{\mathcal{F}}\odot\nabla Q^{\mathcal{F}}$, we deduce that
@@
$-\int_{\mathcal{F}(t)}\partial_{t}Q^{\mathcal{F}}:H\,\mathrm{d}x+\int_{\partial\mathcal{S}(t)}\Bigl(\frac{1}{2}|\nabla Q^{\mathcal{F}}|^{2}+\frac{1}{2}\operatorname{tr}\bigl((Q^{\mathcal{F}})^{2}\bigr)-\frac{1}{3}\operatorname{tr}\bigl((Q^{\mathcal{F}})^{3}\bigr)+\frac{1}{4}\operatorname{tr}\bigl((Q^{\mathcal{F}})^{2}\bigr)^{2}\Bigr)(u^{\mathcal{F}}\cdot\nu(t))\,\mathrm{d}\Gamma.$
@@
$\int_{\partial\mathcal{S}(t)}\Bigl(\frac{1}{2}|\nabla Q^{\mathcal{F}}|^{2}+\frac{1}{2}\operatorname{tr}\bigl((Q^{\mathcal{F}})^{2}\bigr)-\frac{1}{3}\operatorname{tr}\bigl((Q^{\mathcal{F}})^{3}\bigr)+\frac{1}{4}\operatorname{tr}\bigl((Q^{\mathcal{F}})^{2}\bigr)^{2}\Bigr)(u^{\mathcal{F}}\cdot\nu(t))\,\mathrm{d}\Gamma.$
For verifying (a), it thus remains to show that the last three terms in (3.7) cancel out. To this end, we address $\int_{\mathcal{F}(t)}(u^{\mathcal{F}}\cdot\nabla)Q^{\mathcal{F}}:H\,\mathrm{d}x$ and discuss the resulting integrals for each term in $H$. Integrating by parts, and employing $\operatorname{div}u^{\mathcal{F}}=0$, (2.8) and the identities $(u^{\mathcal{F}}\cdot\nabla)((Q^{\mathcal{F}})^{2})=(u^{\mathcal{F}}\cdot\nabla)Q^{\mathcal{F}}\cdot Q^{\mathcal{F}}+Q^{\mathcal{F}}\cdot(u^{\mathcal{F}}\cdot\nabla)Q^{\mathcal{F}}$, $\int_{\mathcal{F}(t)}(u^{\mathcal{F}}\cdot\nabla)Q^{\mathcal{F}}:(\operatorname{tr}((Q^{\mathcal{F}})^{2})\nicefrac{{\operatorname{Id}_{3}}}{{3}})\,\mathrm{d}x=0$ thanks to $\operatorname{tr}(Q^{\mathcal{F}})=0$ as well as the identity involving the above inner product $Q^{\mathcal{F}}:\bigl((u^{\mathcal{F}}\cdot\nabla)(\operatorname{tr}((Q^{\mathcal{F}})^{2})Q^{\mathcal{F}})\bigr)=3(u^{\mathcal{F}}\cdot\nabla)Q^{\mathcal{F}}:\bigl(\operatorname{tr}((Q^{\mathcal{F}})^{2})Q^{\mathcal{F}}\bigr)$, we find that
@@
$=-\int_{\partial\mathcal{S}(t)}\Bigl(\frac{1}{2}\operatorname{tr}\bigl((Q^{\mathcal{F}})^{2}\bigr)-\frac{1}{3}\operatorname{tr}\bigl((Q^{\mathcal{F}})^{3}\bigr)+\frac{1}{4}\operatorname{tr}\bigl((Q^{\mathcal{F}})^{2}\bigr)^{2}\Bigr)(u^{\mathcal{F}}\cdot\nu(t))\,\mathrm{d}\Gamma.$
@@
For the second term, we obtain $\sum_{i,j,k=1}^{3}\int_{\mathcal{F}(t)}u_{k}^{\mathcal{F}}\partial_{k}\nabla Q_{ij}^{\mathcal{F}}\cdot\nabla Q_{ji}^{\mathcal{F}}\,\mathrm{d}x=-\frac{1}{2}\int_{\partial\mathcal{S}(t)}|\nabla Q^{\mathcal{F}}|^{2}(u^{\mathcal{F}}\cdot\nu(t))\,\mathrm{d}\Gamma$, so
(3.9) $\int_{\mathcal{F}(t)}(u^{\mathcal{F}}\cdot\nabla)Q^{\mathcal{F}}:\Delta Q^{\mathcal{F}}\,\mathrm{d}x=-\int_{\mathcal{F}(t)}\nabla u^{\mathcal{F}}:(\nabla Q^{\mathcal{F}}\odot\nabla Q^{\mathcal{F}})\,\mathrm{d}x-\frac{1}{2}\int_{\partial\mathcal{S}(t)}|\nabla Q^{\mathcal{F}}|^{2}(u^{\mathcal{F}}\cdot\nu(t))\,\mathrm{d}\Gamma.$
@@
Concerning (b), we argue that the equilibria $\mathcal{E}$ are the critical points of the energy functional $\mathrm{E}$, so assume $\left.\frac{\,\mathrm{d}}{\,\mathrm{d}t}\mathrm{E}(t)\right|_{t=t_{0}}=0$ at $t_{0}>0$. From (a), it follows that $\nabla u^{\mathcal{F}}=0$ and thus also $u^{\mathcal{F}}=0$ in $\mathcal{F}(t_{0})$ thanks to $u^{\mathcal{F}}=0$ on $\partial\mathcal{O}$. As in the proof of [37, Lemma 2.1], we deduce that $h^{\prime}=\Omega=0$. The relation in (a) also shows that $H(Q_{*})=0$ has to hold for an equilibrium, so $\mathcal{E}$ is of the asserted shape. ∎
Remark $3.$
Combining Section 3(b) and [22, Remark 3.3], stating that $H(Q_{*})=0$ implies $Q_{*}=0$ for spatially constant $Q_{*}$, we conclude that $(0,0,0,0)$ is the only such equilibrium to (2.5)–(2.8).